OpenStudy (anonymous):

Determine the oxidation number of the element that is underlined within the compound. Li2 SiO3 A.+2 B.+4 C.-2 D.-4

OpenStudy (misssunshinexxoxo):

2(+1) + Si + 3(-2) = 0 2 + Si - 6 =0 Si = 6-2 Si = +4